# Free Doubly-Infinitary Distributive Categories are Cartesian Closed

### v6

@@ Abstract.
A central question in category theory concerns the interplay between limits and colimits. The simplest and most structured forms of this interaction are provided by (pseudo)distributivity. Examples include finitary and infinitary distributive categories [10], as well as completely distributive categories [38].
Explicitly, if $\mathbb{C}$ is a category with finite products and finite coproducts, $\mathbb{C}$ is called finitary distributive if the canonical comparison (0.2) induced by the universal property of the coproducts is invertible for any triple $\left(A,B,C\right)$ of objects in $\mathbb{C}$; in other words, $\mathbb{C}$ is finitary distributive if, for any object $A$, the functor
@@
preserves finite coproducts. In the presence of arbitrary coproducts, we can require even more; namely, a category is infinitary distributive if, for any object $A$ and any family $\left(B_{i}\right)_{i\in I}$ of objects, the canonical comparison (0.3) is invertible, which, again, is equivalent to saying that (0.1) is a coproduct-preserving functor for any $A$.
@@
From the perspective of two-dimensional monad theory, both notions can be realized as pseudoalgebras of suitable composites of free completion pseudomonads, arising from canonical (pseudo)distributive laws. By making use of this approach, a much stronger distributivity condition has been considered in [38]; namely, completely distributive categories, that is to say, categories with the distributivity property of arbitrary (small) limits over colimits.
@@
In this paper, we investigate categories equipped with both products and coproducts that satisfy a (pseudo)distributive law, which we call doubly-infinitary distributive categories. This concept serves as an intermediary between infinitary distributive categories and completely distributive ones.
@@
Doubly-infinitary distributive categories naturally appear as pseudoalgebras of a composite pseudomonad, termed herein the free doubly-infinitary distributive category pseudomonad $\mathbf{Dist}$, that arises from a canonical pseudodistributive law between the pseudomonads for free product and free coproduct completions. This pseudodistributive law generalizes the corresponding classical (pseudo)distributive laws between the finite completion counterparts, and constitutes a restriction of the more general pseudodistributive law utilized by Marmolejo et al. [38] to define completely distributive categories.
@@
Since we are dealing with pseudodistributive laws involving lax idempotent pseudomonads, the corresponding composite pseudomonad $\mathbf{Dist}$ is also property-like in the sense of [23]. Furthermore, the $\mathbf{Dist}$-pseudoalgebra structures can be characterized by specifying two properties (corresponding to the pseudoalgebra structures of the original pseudomonads) along with the invertibility of a canonical morphism (a third property, which encompasses the (pseudo)distributivity), e.g., [31, 36, 49].
This situation translates to the following in our context: there is a pseudodistributive law between the lax idempotent pseudomonads for the free completion under products and for the free completion under coproducts. We call the pseudoalgebras of the corresponding composite pseudomonad doubly-infinitary distributive categories. By the above, we conclude that: a category $\mathbb{C}$ is doubly-infinitary distributive if: (1) it has products, (2) it has coproducts, and (3) by denoting $\mathbf{Fam}(\mathbb{C})$ the free coproduct completion of $\mathbb{C}$ (see Section 1 to recall the precise definition), a category $\mathbb{C}$ with arbitrary products and coproducts is doubly-infinitary distributive precisely when the coproduct functor
(0.4) $\coprod:\mathbf{Fam}(\mathbb{C})\to\mathbb{C},$
@@
which endows $\mathbb{C}$ with a coproduct-complete structure (as a pseudoalgebra for the free coproduct completion pseudomonad), preserves products (that is to say, it is a morphism in the ($2$-)category of categories with products, product-preserving functors, and natural transformations).
@@
We explore examples of doubly-infinitary distributive categories, providing a detailed comparative analysis with established categorical properties such as extensivity, infinitary distributivity, and cartesian closedness. Through our investigation, we show that doubly-infinitary distributivity is significantly stricter than classical infinitary distributivity. Notably, we demonstrate that cartesian closedness alone does not imply doubly-infinitary distributivity, even in the presence of arbitrary products and coproducts. Additionally, we identify categories that, while not extensive, still satisfy the conditions of doubly-infinitary distributivity.
@@
Interestingly, even categories that are infinitary extensive, such as the category of topological spaces, fail to be doubly-infinitary distributive. This insight motivates further investigation into identifying precise conditions under which categories of generalized multicategories – for instance, enriched categorical structures and $(T,\mathbb{V})$-categories – exhibit doubly-infinitary distributivity, a topic we leave for future work, see Sect. 5.
@@
Moreover, we illustrate that doubly-infinitary distributivity does not necessarily entail cartesian closedness; for instance, the category of locally connected topological spaces serves as a concrete example of this phenomenon. Nevertheless, we prove that free doubly-infinitary distributive categories are cartesian closed, a property with significant implications for our future work on the denotational semantics of program transformations and related categorical frameworks.
@@
In Section 1, we recall the notion of free (co)product completions. Interested readers will find comprehensive treatments of free (co)product completions $\mathbf{Fam}\left(\mathbb{C}\right)$ of a category $\mathbb{C}$ extensively documented in the literature, notably in [8, 3, 33] and [43, Section 8.5].
Section 3 introduces the pseudomonad structure on free doubly-infinitary distributive categories, denoted by $\mathbf{Dist}$. This structure arises naturally from the canonical pseudodistributive law between the free product completion pseudomad and the free coproduct pseudomonad. We further explore the $2$-category of $\mathbf{Dist}$-pseudoalgebras and pseudomorphisms, whose objects define the class of categories we term doubly-infinitary distributive categories.
@@
Section 4 provides concrete examples of doubly-infinitary distributive categories, contrasting them with established notions including extensivity, cartesian closedness, and distributivity. For additional context on these concepts, readers can consult [8, 10] and especially [43, Section 7].
@@
We note that Von Glehn considered a similar notion of distributivity for $\Pi$- and $\Sigma$-types in fibrations over a locally cartesian closed category $\mathbb{B}$, e.g., [48, 39]. We focus on the specific case of fibrations of the form $\mathbf{Fam}\left(\mathbb{C}\right)\to\mathbf{Set}$, in which case $\Pi$- and $\Sigma$-types reduce to products and coproducts in $\mathbb{C}$ (see [47, Theorem 3.5.2] or [46, Theorem 12]).
@@ 1. Free completion under coproducts
which acts by precomposition with $j(u)$.
The Grothendieck construction applied to this indexed category yields a split fibration $\mathbf{Fam}(\mathbb{C})\to\mathbf{Set}$, whose total category is denoted by $\int\mathfrak{Fam}(\mathbb{C})\cong\mathbf{Fam}(\mathbb{C})$ [19]. The category $\mathbf{Fam}(\mathbb{C})$ admits the following explicit description:
@@
1.1. Universal Property of the $\mathbf{Fam}$-Construction
The assignment $\mathbb{C}\mapsto\mathbf{Fam}(\mathbb{C})$ extends to a lax-idempotent pseudomonad (also called a Kock–Zöberlein pseudomonad [24, 44, 34]) on the $2$-category $\mathbf{Cat}$ of categories, functors, and natural transformations.
@@ 1.1. Universal Property of the $\mathbf{Fam}$-Construction
where $\left(C_{j}\right)_{j\in J}$ is a family of functors $\left(C_{j}:\mathfrak{d}\left(I_{j}\right)\to\mathbb{C}\right)_{j\in J}$, to the object
@@
$\left(f,\left(\phi_{j},\varphi_{j}\right)_{j\in J}\right):\left(J,\left(I_{j},C_{j}\right)_{j\in J}\right)\to\left(J^{\prime},\left(I_{j}^{\prime},C_{j}^{\prime}\right)_{j\in J^{\prime}}\right)$
[…]
$\left(\overline{f},\alpha\right):\left(\coprod_{j\in J}I_{j},\left[\left(C_{j}\right)_{j\in J}\right]\right)\to\left(\coprod_{j\in J^{\prime}}I_{j}^{\prime},\left[\left(C_{j^{\prime}}^{\prime}\right)_{j^{\prime}\in J^{\prime}}\right]\right)$
@@
  • $\overline{f}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\left[\left(\iota_{{}_{I^{\prime}_{f(j)}}}\circ\phi_{j}\right)_{j\in J}\right]$, with $\iota_{{}_{I^{\prime}_{f(j)}}}:I^{\prime}_{f(j)}\to\displaystyle\coprod_{j^{\prime}\in J^{\prime}}I^{\prime}_{j^{\prime}}$ being the coprojections;
  • by observing that $\left[\left(C_{j^{\prime}}^{\prime}\right)_{j^{\prime}\in J^{\prime}}\right]\circ\overline{f}=\left[\left(C^{\prime}_{f(j)}\circ\phi_{j}\right)_{j\in J}\right]$, we define
    $\alpha\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\left[\left(\varphi_{j}:C_{j}\to C^{\prime}_{j}\circ\phi_{j}\right)_{j\in J}\right],$
@@
Moreover, the $2$-category of pseudoalgebras and pseudomorphisms for this pseudomonad is equivalent to $\mathbf{CoProdCat}$, whose:
@@
  • $1$-cells are functors preserving these coproducts, and
  • $2$-cells are natural transformations.
It is well known that $\mathbf{Fam}$ is a pseudomonad that satisfies the Kock–Zöberlein condition, that is to say, it is a lax idempotent pseudmonad. This can be more easily verified by verifying that the pseudomonadic biadjunction
@@
satisfies the following: we have an adjunction $G\epsilon\dashv\eta G$, and it actually forms a lali (left adjoint left inverse); in other words, indeed, for any coproduct-complete category $\mathbb{C}$,
@@
has a right adjoint which is precisely $\left(G\epsilon\right)_{\mathbb{C}}$. For background on such lali adjunctions, see [27, Definition 2.5], and for the Kock–Zöberlein condition in this context, see [13, Theorem 3.15].
One obtains a colax-idempotent pseudomonad on $\mathbf{Cat}$ via the composite
@@
whose pseudoalgebras correspond to categories equipped with products. The resulting $2$-category of categories with products, products-preserving functors and natural transformations is denoted by $\mathbf{ProdCat}$.
@@
(1.1) $\langle C\mid I\rangle=\langle C(i)\mid i\in I\rangle=\langle C_{i}\mid i\in I\rangle$
to suggest their interpretation as freely generated products of a family $(C_{i})_{i\in I}$ of objects in $\mathbb{C}$, while we denote by
(1.2) $\left[C\mid I\right]=\left[C(i)\mid i\in I\right]=\left[C_{i}\mid i\in I\right]$
the corresponding object in $\mathbf{Fam}(\mathbb{C})$, again to emphasize their interpretation as freely generated coproducts.
@@ 2. The Free Doubly-Infinitary Distributive Category
In this section, given a category $\mathbb{C}$, we explicitly define the category $\mathbf{Dist}\left(\mathbb{C}\right)$, and establish our main result on its cartesian closedness, presented in Theorem 2.4. We call the category $\mathbf{Dist}\left(\mathbb{C}\right)$ the free doubly-infinitary distributive category on $\mathbb{C}$. This terminology is further justified in Section 3, where we establish that $\mathbf{Dist}$ is actually part of a pseudomonad whose pseudoalgebras are the doubly-infinitary distributive categories, as defined in 3.1.
@@
The universal property of (2.1) will be formalized in Section 3. By construction, $\mathbf{Dist}\left(\mathbb{C}\right)$ admits coproducts, being a free coproduct completion. In addition, it admits products and, more unexpectedly, exponentials. We now provide an explicit description of $\mathbf{Dist}\left(\mathbb{C}\right)$, which will make these structures fully transparent.
Explicit description of the $\mathbf{Dist}$-construction.
@@
$\left[\langle C(j,i)\mid i\in I_{j}\rangle\mid j\in J\right]$
@@
A morphism $\left[\langle C(j,i)\mid i\in I_{j}\rangle\mid j\in J\right]\to\left[\langle C^{\prime}({j^{\prime},i^{\prime}})\mid i^{\prime}\in I^{\prime}_{j^{\prime}}\rangle\mid j^{\prime}\in J^{\prime}\right]$ consists of a triple
(2.2) $\left(g,f,\alpha\right):\left[\langle C(j,i)\mid i\in I_{j}\rangle\mid j\in J\right]\to\left[\langle C^{\prime}(j^{\prime},i^{\prime})\mid i^{\prime}\in I_{j^{\prime}}^{\prime}\rangle\mid j^{\prime}\in J^{\prime}\right]$
where $g:J\to J^{\prime}$ is a function, and $f=\left(f_{j}:I_{g(j)}^{\prime}\to I_{j}\right)_{j\in J}$ is a family of functions, and
@@
$\displaystyle\mathbf{Dist}\left(\mathbb{C}\right)\left(\left[\langle C(j,i)\mid i\in I_{j}\rangle\mid j\in J\right],\left[\langle C^{\prime}(j^{\prime},i^{\prime})\mid i^{\prime}\in I^{\prime}_{j^{\prime}}\rangle\mid j^{\prime}\in J^{\prime}\right]\right)$
=$\stackrel{{\scriptstyle\mathrm{def}}}{{=}}$
@@
The identity morphism on an object $\left[\langle C\left(j,i\right)\mid i\in I_{j}\rangle\mid j\in J\right]$ is given by
@@
(2.4) $\left(e,d,\beta\right):\left[\langle C^{\prime}(j^{\prime},i^{\prime})\mid i^{\prime}\in I_{j}^{\prime}\rangle\mid j^{\prime}\in J^{\prime}\right]\to\left[\langle C^{\prime\prime}(j^{\prime\prime},i^{\prime\prime})\mid i^{\prime\prime}\in I_{j}^{\prime\prime}\rangle\mid j^{\prime\prime}\in J^{\prime\prime}\right]$
@@
Products and coproducts in $\mathbf{Dist}\left({\mathbb{C}}\right)$

### v7

@@ Abstract.
A central question in category theory concerns the interplay between limits and colimits. The simplest and most structured forms of this interaction are captured by (pseudo)distributivity. Prominent examples include finitary and infinitary distributive categories [10] and completely distributive categories [42].
Explicitly, if $\mathbb{C}$ is a category with finite products and finite coproducts, then $\mathbb{C}$ is called finitary distributive if the canonical comparison (0.2), induced by the universal property of coproducts, is invertible for any triple $\left(A,B,C\right)$ of objects in $\mathbb{C}$. Equivalently, $\mathbb{C}$ is finitary distributive if, for every object $A$, the functor
@@
preserves finite coproducts. If arbitrary coproducts exist, one can strengthen this requirement by demanding that, for every object $A$ and every family $\left(B_{i}\right)_{i\in I}$ of objects, the canonical comparison (0.3) is invertible. This condition is equivalent to the preservation of all coproducts by the functor (0.1).
@@
From the viewpoint of two-dimensional monad theory, both notions can be realized as pseudoalgebras for suitable composites of free completion pseudomonads, arising from canonical (pseudo)distributive laws. Within this approach, a stronger distributivity condition was introduced in [42], namely that of completely distributive categories, in which arbitrary (small) limits distribute over colimits.
@@
In this paper, we investigate categories equipped with both products and coproducts that satisfy a (pseudo)distributive law, which we call doubly-infinitary distributive categories. This concept serves as an intermediate notion between infinitary distributive and completely distributive categories.
@@
Doubly-infinitary distributive categories arise as pseudoalgebras for the composite pseudomonad that we call the free doubly-infinitary distributive category pseudomonad $\mathbf{Dist}$. This pseudomonad results from the canonical pseudodistributive law between the pseudomonads for free product and free coproduct completions. It generalizes the corresponding classical finite (pseudo)distributive laws and constitutes a restriction of the pseudodistributive law used by Marmolejo et al. [42] to define completely distributive categories.
@@
We work with pseudodistributive laws between Kock–Zöberlein pseudomonads (also known as lax idempotent pseudomonads[27, 38, 14]. In this context, pseudoalgebra structures can be characterized by the satisfaction of two properties corresponding to the underlying pseudomonads, together with the invertibility of a canonical morphism expressing the (pseudo)distributivity condition [34, 40, 56, 36]. Consequently, the resulting composite pseudomonad is property-like in the sense of [26].
Explicitly, we show that a category $\mathbb{C}$ is doubly-infinitary distributive if (1) it has products, (2) it has coproducts, and (3) for every family of families of objects $\bigl(\bigl(C_{ji}\bigr)_{i\in I_{j}}\bigr)_{j\in J}$ in $\mathbb{C}$, the canonical morphism (3.1) is invertible. Equivalently, if $\mathbf{Fam}(\mathbb{C})$ denotes the free coproduct completion of $\mathbb{C}$ (see Section 1 for the precise definition), then a category $\mathbb{C}$ with arbitrary products and coproducts is doubly-infinitary distributive precisely when the coproduct functor
(0.4) $\coprod:\mathbf{Fam}(\mathbb{C})\to\mathbb{C}$
@@
preserves products, which means that it is a morphism in the $2$-category of categories with products and product-preserving functors.
@@
We explore examples of doubly-infinitary distributive categories and provide a detailed comparative analysis with established categorical properties such as extensivity, infinitary distributivity, and cartesian closedness. Our investigation shows that doubly-infinitary distributivity is substantially more restrictive than its classical infinitary counterpart. Even categories that are both cocomplete and infinitary lextensive, such as the category of topological spaces, fail to be doubly-infinitary distributive.
@@
These observations motivate further investigation into the precise conditions under which categories of generalized multicategories, such as enriched categorical structures, $(T,\mathbb{V})$-categories, and their associated graph-like notions, exhibit doubly-infinitary distributivity. We leave this topic for future work; see Section 6.
Furthermore, we show that cartesian closedness, even when combined with completeness and cocompleteness, does not entail doubly-infinitary distributivity. For instance, the category of pseudo-topological spaces is a complete and cocomplete quasitopos but is not doubly-infinitary distributive.
@@
Finally, we show that doubly-infinitary distributivity does not in general imply cartesian closedness. The category of locally connected topological spaces provides a concrete example of this phenomenon. Nevertheless, we prove that free doubly-infinitary distributive categories are cartesian closed, a property that has important implications for our future work on the denotational semantics of program transformations and related categorical frameworks.
@@
In Section 1, we recall the notion of free (co)product completions. Interested readers will find comprehensive treatments of free (co)product completions $\mathbf{Fam}\left(\mathbb{C}\right)$ of a category $\mathbb{C}$ extensively documented in the literature, notably in [8, 3, 37] and [48, Section 8.5].
Section 3 introduces the pseudomonad structure on free doubly-infinitary distributive categories, denoted by $\mathbf{Dist}$. This structure arises naturally from the canonical pseudodistributive law between the free product completion pseudomonad and the free coproduct pseudomonad. We further explore the $2$-category of $\mathbf{Dist}$-pseudoalgebras and pseudomorphisms, whose objects define the class of categories we term doubly-infinitary distributive categories.
@@
Section 5 provides concrete examples of doubly-infinitary distributive categories, contrasting them with established notions including extensivity, cartesian closedness, and distributivity. For additional context on these concepts, readers can consult [8, 10] and especially [48, Section 7].
@@
We note that Von Glehn considered a similar notion of distributivity for $\Pi$- and $\Sigma$-types in fibrations over a locally cartesian closed category $\mathbb{B}$, e.g., [55, 43]. We focus on the specific case of fibrations of the form $\mathbf{Fam}\left(\mathbb{C}\right)\to\mathbf{Set}$, in which case $\Pi$- and $\Sigma$-types reduce to products and coproducts in $\mathbb{C}$ (see [54, Theorem 3.5.2] or [53, Theorem 12]).
@@ 1. Free completion under coproducts
which acts by precomposition with $\mathfrak{d}(u)$.
The Grothendieck construction applied to this indexed category yields a split fibration $\mathbf{Fam}(\mathbb{C})\to\mathbf{Set}$, whose total category is denoted by $\int\mathfrak{Fam}(\mathbb{C})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\Sigma_{\mathbf{Set}}\mathfrak{Fam}(\mathbb{C})\cong\mathbf{Fam}(\mathbb{C})$, e.g. [21] and [37, Section 6]. The category $\mathbf{Fam}(\mathbb{C})$ admits the following explicit description:
@@
Universal Property of the $\mathbf{Fam}$-Construction
The assignment $\mathbb{C}\mapsto\mathbf{Fam}(\mathbb{C})$ extends to a lax-idempotent pseudomonad (also called a Kock–Zöberlein pseudomonad, e.g. [27, 38, 49, 14]) on the $2$-category $\mathbf{Cat}$ of categories, functors, and natural transformations.
@@
where $\left(I_{j},C_{j}\right)_{j\in J}$ is a family of functors $\left(C_{j}:\mathfrak{d}\left(I_{j}\right)\to\mathbb{C}\right)_{j\in J}$, to the object
@@
$\left(f,\left(h_{j},\varphi_{j}\right)_{j\in J}\right):\left(J,\left(I_{j},C_{j}\right)_{j\in J}\right)\to\left(J^{\prime},\left(I_{j^{\prime}}^{\prime},C_{j^{\prime}}^{\prime}\right)_{j^{\prime}\in J^{\prime}}\right)$
in $\mathbf{Fam}(\mathbf{Fam}(\mathbb{C}))$, where each $\left(h_{j},\varphi_{j}\right):\left(I_{j},C_{j}\right)\to\left(I^{\prime}_{f(j)},C^{\prime}_{f(j)}\right)$ is a morphism in $\mathbf{Fam}(\mathbb{C})$, to the morphism in $\mathbf{Fam}(\mathbb{C})$ defined by
$\left(\overline{f},\alpha\right):\left(\coprod_{j\in J}I_{j},\left[\left(C_{j}\right)_{j\in J}\right]\right)\to\left(\coprod_{j^{\prime}\in J^{\prime}}I_{j}^{\prime},\left[\left(C_{j^{\prime}}^{\prime}\right)_{j^{\prime}\in J^{\prime}}\right]\right)$
@@
  • $\overline{f}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\left[\left(\iota_{{}_{I^{\prime}_{f(j)}}}\circ h_{j}\right)_{j\in J}\right]$, with $\iota_{{}_{I^{\prime}_{f(j)}}}:I^{\prime}_{f(j)}\to\displaystyle\coprod_{j^{\prime}\in J^{\prime}}I^{\prime}_{j^{\prime}}$ being the coprojections;
  • by observing that $\left[\left(C_{j^{\prime}}^{\prime}\right)_{j^{\prime}\in J^{\prime}}\right]\circ\overline{f}=\left[\left(C^{\prime}_{f(j)}\circ h_{j}\right)_{j\in J}\right]$, we define
    $\alpha\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\left[(\varphi_{j}:C_{j}\to C^{\prime}_{f(j)}\circ h_{j})_{j\in J}\right],$
@@
The $2$-category of pseudoalgebras and pseudomorphisms for this pseudomonad is equivalent to $\mathbf{CoProdCat}$, whose:
@@
  • 11-cells are functors preserving these coproducts, and
  • 22-cells are natural transformations.
It is well known that $\mathbf{Fam}$ is a pseudomonad that satisfies the Kock–Zöberlein condition, that is to say, it is a lax idempotent pseudomonad. This can be more easily verified by showing that the pseudomonadic biadjunction
@@
satisfies the following: we have a lali (left adjoint left inverse) adjunction $G\epsilon\dashv\eta G$; in other words, indeed, for any coproduct-complete category $\mathbb{C}$,
@@
has a left adjoint which is precisely $\left(G\epsilon\right)_{\mathbb{C}}$. For background on such lali adjunctions, see [30, Definition 2.5], and for the Kock–Zöberlein condition in this context, see [14, Theorem 3.15].
By duality, one obtains a colax-idempotent pseudomonad on $\mathbf{Cat}$ via the composite
@@
whose pseudoalgebras correspond to categories equipped with products. The resulting $2$-category of categories with products, product-preserving functors and natural transformations is denoted by $\mathbf{ProdCat}$.
@@
(1.1) $\langle C|I\rangle=\langle C(i)|i\in I\rangle=\langle C_{i}|i\in I\rangle$
to suggest their interpretation as freely generated/formal products of a family $(C_{i})_{i\in I}$ of objects in $\mathbb{C}$, while we denote by
(1.2) $\left[C|I\right]=\left[C(i)|i\in I\right]=\left[C_{i}|i\in I\right]$
the corresponding object in $\mathbf{Fam}(\mathbb{C})$, again to emphasize their interpretation as formal coproducts.
@@ 2. The Free Doubly-Infinitary Distributive Category
In this section, given a category $\mathbb{C}$, we explicitly define the category $\mathbf{Dist}\left(\mathbb{C}\right)$, and establish our main result on its cartesian closedness, presented in Theorem 2.4. We call the category $\mathbf{Dist}\left(\mathbb{C}\right)$ the free doubly-infinitary distributive category on $\mathbb{C}$. This terminology is further justified in Section 3, where we establish that $\mathbf{Dist}$ is actually part of a pseudomonad whose pseudoalgebras are the doubly-infinitary distributive categories, as defined in Def. 3.1.
@@
The universal property of (2.1) will be formalized in Section 3. By construction, $\mathbf{Dist}\left(\mathbb{C}\right)$ admits coproducts, being a free coproduct completion. In addition, it admits products and, perhaps unexpectedly, exponentials. We now provide an explicit description of $\mathbf{Dist}\left(\mathbb{C}\right)$, which will make these structures fully transparent.
2.1. Explicit description of the $\mathbf{Dist}$-construction
@@
$\left[\langle C(j,i)|i\in I_{j}\rangle|j\in J\right]$
@@
A morphism $\left[\langle C(j,i)|i\in I_{j}\rangle|j\in J\right]\to\left[\langle C^{\prime}({j^{\prime},i^{\prime}})|i^{\prime}\in I^{\prime}_{j^{\prime}}\rangle|j^{\prime}\in J^{\prime}\right]$ consists of a triple
(2.2) $\left(g,f,\alpha\right):\left[\langle C(j,i)|i\in I_{j}\rangle|j\in J\right]\to\left[\langle C^{\prime}(j^{\prime},i^{\prime})|i^{\prime}\in I_{j^{\prime}}^{\prime}\rangle|j^{\prime}\in J^{\prime}\right]$
where $g:J\to J^{\prime}$ is a function, $f=\left(f_{j}:I_{g(j)}^{\prime}\to I_{j}\right)_{j\in J}$ is a family of functions, and
@@
$\displaystyle\mathbf{Dist}\left(\mathbb{C}\right)\left(\left[\langle C(j,i)|i\in I_{j}\rangle|j\in J\right],\left[\langle C^{\prime}(j^{\prime},i^{\prime})|i^{\prime}\in I^{\prime}_{j^{\prime}}\rangle|j^{\prime}\in J^{\prime}\right]\right)$
$\stackrel{{\scriptstyle\mathrm{def}}}{{=}}$
@@
The identity morphism on an object $\left[\langle C\left(j,i\right)|i\in I_{j}\rangle|j\in J\right]$ is given by
@@
(2.4) $\left(e,d,\beta\right):\left[\langle C^{\prime}(j^{\prime},i^{\prime})|i^{\prime}\in I_{j^{\prime}}^{\prime}\rangle|j^{\prime}\in J^{\prime}\right]\to\left[\langle C^{\prime\prime}(j^{\prime\prime},i^{\prime\prime})|i^{\prime\prime}\in I_{j}^{\prime\prime}\rangle|j^{\prime\prime}\in J^{\prime\prime}\right]$
@@
2.2. Products and coproducts in $\mathbf{Dist}\left({\mathbb{C}}\right)$